The probability of a Hat being red is 1/4.
The probability of a Hat being green is 1/4
The probability of a Hat being white is 1/2
The probability of a Hat being red or white is __

Answer by Edwin McCravy(20054)

P(R or W) = P(R) + P(W) - P(R and W)

The hat can't be both red and white because 

P(R) + P(G) + P(W) = 1/4 +  1/4 + 1/2 = 1, so P(R and W) = 0

P(R or W) = P(R) + P(W) - P(R and W) = 1/4 + 1/2 - 0 = 1/4 + 2/4 = 3/4 

Or we could have just done it this way: 

P(R or W) = P(not G) = 1 - 1/4 = 4/4 - 1/4 = 3/4

Edwin